OpenStudy (anonymous):

For which acute angle are the sine of the angle and cosine of the angle equal? 30º 45º 60º <--- My guess. 75º

OpenStudy (anonymous):

\[\sin \theta=\cos \theta \] \[\frac{ \sin \theta }{ \cos \theta }=1\] \[\tan \theta=1\] \[\theta=? \]

OpenStudy (anonymous):

your guess is not correct.

OpenStudy (anonymous):

45